isanaiscool:

Sorry ill rewrite the question (Technical difficulties) an airplane is flying from montreal to vancouver. the wind is blowing from the west at 60km/h. the airplane flies at an airspeed of 750km/h and must stay on a heading of 65 degrees west of north. what heading should the pilot take to compensate for the wind? what is the speed of the airplane relative to the ground?

Narad:

Make a sketch Let angle of flight be N(a+65)ºW and speed be d mph Applying sine rule \[\frac{ 60 }{ sina }=\frac{ 750 }{ \sin155º}\] \[\frac{ d }{ \sin(180-155-a) }=\frac{ 750 }{ sin155º }\] Solving those 2 equations, you get the direction of flight and the relative speed of the plane relative to the ground